# Influencer Identification on Link Predicted Graphs

### v2

@@ Abstract
link prediction, centrality metrics, vital nodes identification
I Introduction
@@
When a company is launching a new product and wants to use social media for marketing, understanding who to hire for such campaigns becomes of most importance to maximize profit. With the explosion of social media in recent years, a common strategy has been to target certain influential individuals for advertising. These individuals, or “influencers,” become the initial spreaders in different social settings, and through their connections and followers, companies reach greater awareness (and sales) in the entire society. However, social groups are constantly evolving and thus influencers in certain social setting may no longer be the optimal selection in the few months or years until the product is released, which can lower efficiency. It becomes clear, then, that being able to predict future influencers is extremely important.
@@
Within a mathematical setting, one can consider social networks as graphs, where the nodes represent members of the network, and their connections are represented by edges. The study of how nodes can influence a network has been done for many decades, and the reader can refer to recent surveys such as Lü et al.’s listofidentificationtypes in 2016, Pei et al.’s survey2 in 2019, and AbdulAmeer et al.’s survey1 in 2022 to learn about the scope of such work.
@@
In particular, Domingos and Richardson imp1 studied influence through a viral marketing perspective. In 2003, Kempe et al. imp2 formalized this as the influence maximization problem (IMP) of choosing $k$ nodes of greatest influence, proving that a complete solution was NP-hard. However, it is interesting to note that to the best of our knowledge, no mathematical definition of a social Influencer as been proposed in the literature, and thus we shall introduce here what we think can be a useful way of thinking of such important individuals within our society (see Definition 29).
@@
In the present paper we will focus on heuristic measures, revolving around local centralities due to lower time complexity: while global metrics take into consideration more complete information about the graph, they are often computationally expensive. Furthermore, Hu et al. hu claimed that global influence can be approximated by local influence from neighbors of order around three to four, which means local measures may be more cost-effective. As mentioned before, one should keep in mind that social networks are constantly changing. In 2004, Liben-Nowell and Kleinberg linkpredictioncomparison formalized the link prediction problem for social networks as predicting future edges based on knowledge of the current graph. Hasan and Zaki linkpredictionsurvey and Lü and Zhou linkpredictionsurvey2 provided surveys of link prediction results. In our work, we shall focus on local similarity metrics and their quasi-local extensions given by Aziz et al. globalextensions , which are heuristic measures similar to centrality.
@@
In 2019, Ghafouri and Khasteh lpim1 introduced the idea of using a graph modelling technique to engage in link prediction prior to use of greedy algorithms for influence maximization as a way to account from invisible edges between nodes (i.e. edges not observed when collecting data on the network). Singh and Kailasam lpim2 utilized a similar approach when predicting sets of influential nodes for the next iteration of the graph through link prediction using the a type of Restricted Boltzmann Machine (RBM). In 2023, Yanchenko et al. exante discussed the use of deep learning link prediction to understand the outcomes of the future networks, then identify influential nodes through the greedy algorithm and dynamic degree discount. We take a similar approach, but focus on heuristic link prediction measures whereas they consider linear regression, deep learning, matrix factorization, and graph neural networks, and more. We also evaluate many other common centrality metrics and algorithms that, to the best of our knowledge, previous literature has not covered. In an effort to lower time complexity, our graph prediction also makes use of expected value to decrease computation cost of influence spread.
@@
We shall seek to quantify influence through simple and complex contagion. As stated by Min & Miguel simplecontagion , simple contagion is a model from epidemiology that runs on the assumption that there is a fixed probability of transmission between each pair of vertices. Centola and Macy complexcontagiondefinition defined complex contagion as a model for actions or information with some perceived social cost, considering both personal resistance to influence and forces like peer pressure and social affirmation: a person is infected only if sufficiently many of their neighbors are.
@@
In general terms, easily accepted information spreads as simple contagion whereas more controversial or effort-requiring topics may be complex contagion. Examples of some categorizations can be seen in Table 1 below. We shall begin our study in Section II) by studying link prediction and influencer identification separately on dynamic social networks, and then we shall examine influence through simple and complex contagion models in Section IV.2). By putting the above together we shall present our definition of influencer, detail the specific influence model definitions, and introduce our concrete algorithm, the Social Sphere Model (see Section III).
@@
Viral Memes
Simple Contagion
Weng et al. memescontagions
Nonviral Memes
Complex Contagion
Weng et al. memescontagions
Spread of Disease
Simple Contagion
Health Behaviors
Complex Contagion
Campbell and Salath’e healthbehaviors
Music
Simple Contagion
Notarmuzi et al. conversationcontroversial
TV Shows
Simple Contagion
Notarmuzi et al. conversationcontroversial
Political/Societal Controversies
Complex Contagion
Notarmuzi et al. conversationcontroversial
Social Movements
Complex Contagion
Centola change
@@
Researchers have noted that the performances of the same set of nodes differ noticeably in the simple and complex contagion (see Figure 1 for an example). For example, Watts and Dodds ltmbehavior found that while the former places a great deal of emphasis on influencers, the latter says away from choosing traditional influencers as initial nodes, especially when considering the relative costs (traditionally influential nodes are likely more expensive hires).
@@
In Section V, and through the use of the Social Sphere Model, we observe that influence generally spreads much faster in the simple contagion model than in the complex. We also notice the following:
  • better performance for some parameters in our modified prediction metrics and algorithms when compared to the originals (see Section II);
  • high similarity between our model’s predicted influencers and the ‘true’ future influencers in both selection and collective influence, suggesting the effectiveness of our model; and
  • potential in empirical datasets for identifying influencers overlooked by a static model.
Our paper is organized as follows: in Section II, we explore previous works around heuristic measures in link prediction and influencer identification. We explain the Social Sphere Model and our mathematical definition for influencers in Section III. Our methodology is detailed in Section IV, with evaluation methods in Section IV.1, concrete contagion models in Section IV.2, and experimental setup in Section IV.3. We examine our model’s influence in simple and complex contagion through analysis of our algorithm’s results on both empirical datasets and random graphs in Section V. Finally, we give a summary and future directions for our research in Section VI.
II Social Networks
@@
where $w(u,v)$ is the weight of edge $uv$. Such weights may represent a variety of parameters of the social network, and in this paper, we consider weights to be approximations for probabilities of influence transmission, be it information, disease, or more. In particular, we take inspiration from Goyal et al. bernoullidistribution , who took the weights $p(i,j)$ to be the total number of interactions over the total number of trials. In order to study such models, one may assume that the probability of interaction between two nodes during a fixed interval of time is constant, in which case each interval of time can be considered as a Bernoulli trial for temporal datasets.
II.1 Weight functions
@@ II.2 Link Prediction
Common Neighbors. Considering collaboration networks, Newman proposed in commonneighbors that the more common neighbors two nodes share, the more likely they are to form a link in the future, leading to the following definition:
@@
Murata and Moriyasu weightedlinkprediction amended common neighbors for weighted networks to be dependent on the sum of the weights between the nodes summed over common neighbors leading to the following weighted definition:
@@
Local Path (Quasi-Local Common Neighbors). Zhou et al. resourceallocation created the local path metric from the observation that common neighbors as a similarity metric often gives the same scores to many nodes, making the actual rankings ambiguous. They extended the computation of score to include paths of three edges, weighting the longer paths by some $\epsilon$:
@@
Here, we take $\epsilon=10^{-3}$. Moreover, Aziz et al. globalextensions noted that local path is the quasi-local extension of common neighbors. Extending on the previously mentioned work of Murata and Moriyasu weightedlinkprediction , Bai et al. weightedlocalpath developed a weighted index for local path (weights here are adjusted to match the latter):
@@
Jaccard Coefficient. Liben-Nowell and Kleinberg linkpredictioncomparison derived the Jaccard coefficient from information retrieval, with the following definition:
@@
Resource Allocation. Zhou et al. resourceallocation came up with the resource allocation metric based on the eponymous process where each individual has resources to be split evenly among its neighbors, defined as follows:
@@
Lü and Zhou weightednetworkweakties made a variant of the above resource allocation for weighted graphs, leading to the following:
@@
Quasi-Local Resource Allocation. Aziz et al. globalextensions made a quasi-local extension for RA, where one considers paths of three edges in addition to paths of two edges. In other words:
@@
Quasi-Local variant of RA22-2- 2 Inspired by Aziz et al. globalextensions , we also create a quasi-local variant for RA22-2- 2:
@@ II.3 Node Centrality
Degree Centrality
$O(|E|)$degreecomplexity
$O(|E|)$degreecomplexity
local listofinfluencers2
Coreness
$O(|E|)$corenessalgorithm
$O(|E|)$
iterative listofinfluencers2
H-index
$O(|V|\langle k\rangle)$ h-indexcomplexity
$O(|V|\langle k\rangle)$
local listofinfluencers2
LocalRank
$O(|V|\langle k\rangle^{2})$ localrank
N/A
local listofinfluencers2
Clustering Coefficient (and ClusterRank)
$O(|V|\max_{v\in V}k_{v}^{2})$ clusteringcoefftime
$O(|V|\max_{v\in V}k_{v}^{2})$
listofinfluencers2
Closeness
$O(|V||E|)$ timecomplexities
$O(|V||E|+|V|^{2}\log|V|)$ fibheap
path-based listofidentificationtypes , global listofinfluencers2
Betweenness
$O(|V||E|)$ betweennessalgorithm
$O(|V||E|+|V|^{2}\log|V|)$ betweennessalgorithm
path-based listofidentificationtypes , global listofinfluencers2
Eigenvector
$O(|V|^{2}\cdot\text{iteration \#})$ poweriterationmethod
$O(|V|^{2}\cdot\text{iteration \#})$ poweriterationmethod
global, iterative listofinfluencers2
PageRank
$O(|E|\cdot\text{iteration \#})$pagerankcomplexity
$O(|E|\cdot\text{iteration \#})$pagerankcomplexity
iterative listofinfluencers2
Balanced Index
$O(\langle k\rangle V+V\log V)$ ltmmetrics
$O(\langle k\rangle V+V\log V)$
local
@@
Vital node identification is a crucial part of social network analysis, applicable to many field ranging from marketing to epidemiology. Centrality methods, some of the most common of these processes, offer a window into understanding influence dynamics within social networks. Mathematically, centrality methods are based on functions that assign a real-number score to each $v\in V$, essentially ranking the nodes in terms of influence listofidentificationtypes .
@@
Coreness. In 1983, Seidman kcoreorigin defined the $k$-core as a measure of the network cohesion of a node, where a subgraph is a $k$-core if it has minimum degree at least $k$. Batagelj and Zaveršnik corenessalgorithm define a $k$-core decomposition algorithm in Algorithm 1.
Algorithm 1 $k$-shell decomposition corenessalgorithm
@@
Later on, Kitsak et al. coreness took the core numbers for each vertex as its corenness value. Eidsaa and Almaas s-core created a weighted version of $k$-core, where node strength is used rather than degree in the decomposition and when a node $u$ is selected (has the minimum strength in the pruned graph), it is in the $\lceil s_{u}\rceil$-core. A similar definition can be taken using $\lfloor s_{u}\rfloor$, and the algorithm is modified accordingly in Algorithm 2 below.
@@
H-index. Along the same lines of study, but within the academic world, Hirsch h-indexorigin proposed the H-index to measure the impact of a scientist’s research outputs through considering citations per paper. Lü et al. h-index then used it in the context of centrality metrics, introducing the following definition:
@@
From the above perspective, the classical H-index can be seen as the $1$-order H-index. Lü et al. h-index proved that the H-index forms a middle ground between a progression from degree centrality to coreness, where the 00-order index is degree and the $\infty$-order index converges to the $k$-core score. Finally, it is interesting to note that Zhao et al. weightedhindex defined the w-lobby index, a metric for weighted networks similar to the H-index, as follows:
@@
LocalRank. Aiming to find a more effective local method due to the high time complexities of global and path-based methods, Chen et al. localrank introduced the following extension of degree centrality to more than just first-order neighborhoods:
@@
ClusterRank. Watts and Strogatz clusteringcoeff defined the clustering coefficient around the percent of existing links between a node’s neighbors over the total potential amount to quantify how clustered the network is:
@@
Working with the intuition that nodes in more clustered parts of the network may be more redundant and thus have less influence, Chen et al. clusterrank then used the clustering coefficients of nodes to define the following centrality metric:
@@
Closeness. Bavelas introduced in 1948 closenessorigin the idea of closeness centrality as proportional to the reciprocal of the sum of the shortest path lengths from each pair of vertices $u\neq v$. Rephrasing, Freeman closeness defines the concept of closeness centrality as a measure of the relative proximity of a node $p_{k}$ to the rest of the graph $G$, defined as
@@
In a survey of influencer identification metrics, Lü et al. listofidentificationtypes note that disconnected networks then pose a problem as distances go to infinity, making the nodes indistinguishable by closeness centrality (as they are all then 00). They then give the following modification:
@@
Betweenness. Betweenness centrality considers how critical a node $i$ is to the spread of information, e.g. what percentage of shortest paths going from one node to another must pass through $i$. First formally defined it in 1977 by Freeman betweennessorigin , it can be understood along with an efficient algorithm for computing betweenness centrality, through the work of Brandes betweennessalgorithm :
@@
Eigenvector. In eigenvectororigin Landau gave a clarified overview of a previous paper of his detailing the use of eigenvectors to rank chess players, which has been credited with being the first use of eigenvector centrality. Later on, Lü et al. listofidentificationtypes defined the eigenvector centrality scores $EC$ of $G$ as the vector $\vec{x}$ with components $x_{i}=EC_{i}$ such that $\vec{x}:=\frac{1}{\lambda}A\vec{x}$ for the largest eigenvector $\lambda$ of $A$. In other words,
@@
PageRank. PageRank is a search algorithm presented by Brin and Page pagerankorigin in a paper where they introduced a Google prototype, modeled on the behavior of a bored, link-clicking user. They defined the algorithm as follows:
@@
Xing and Ghorbani weightedpagerank modified the PageRank algorithm for weighted graphs:
@@
LeaderRank. Inspired by PageRank, Lü et al. leaderrank created the LeaderRank algorithm. Their intention was to include a “ground node” connected to every node in the network, called node $n+1$, and then used the following recursive relationship to solve for their influences:
@@
Karampourniotis et al. ltmmetrics introduced the Balanced Index metric for complex contagion by taking a weighted combination the metrics of resistance, degree, and neighbors’ degrees. In the context of threshold models, the resistance $r_{v}$ of a node $v$ is defined as a measure of the current threshold of each node, where uninfected nodes $v$ initially begin with $r_{v}=\theta_{v}\cdot s(v)$ and experience decreasing resistance as more of its neighbors become infected. At resistance 00 the node also becomes infected. The neighbors’ degree comes from Karampourniotis et al.’s ltmmetrics indirect drop of resistance metric, which for node $v$ takes the sum over the nodes that $v$ could infect in a single turn (i.e. the nodes $j$ with remaining resistance less than $\frac{w(i,j)}{s(i)}$). The balanced index metric is thus as follows:
@@
Complex Path Centrality. Seeking to extend the path length measures traditionally used in simple contagion to the wide bridges (reinforcing ties) used in complex contagion paths, Guilbeault and Centola complexpathlength created complex path centrality, based on the following:

### v3

# Formal analysis, Investigation, Writing - Review & Editing

@@ Abstract
1 Introduction
@@
When a company is launching a new product and wants to use social media for marketing, understanding who to hire for such campaigns becomes of most importance to maximize profit. With the explosion of social media in recent years, a common strategy has been to target certain influential individuals for advertising. These individuals, or “influencers," become the initial spreaders in different social settings, and through their connections and followers, companies reach greater awareness (and sales) in the entire society. However, social groups are constantly evolving and thus influencers in certain social setting may no longer be the optimal selection in the few months or years until the product is released, which can lower efficiency. As a result, being able to predict future influencers is extremely important.
@@
Within a mathematical setting, one can consider social networks as graphs, where the nodes represent members of the network, and their connections are represented by edges. The study of how nodes can influence a network has been done for many decades, and the reader can refer to recent surveys such as Lü et al.’s [42] in 2016, Pei et al.’s [50] in 2019, and AbdulAmeer et al.’s [1] in 2022.
@@
In particular, Domingos and Richardson [16] studied influence through a viral marketing perspective. In 2003, Kempe et al. [32] formalized this as the influence maximization problem (IMP) of choosing $k$ nodes of greatest influence, proving that a complete solution was NP-hard. However, it is interesting to note that to the best of our knowledge, no mathematical definition of a social Influencer as been proposed in the literature, and thus we shall introduce here what we think can be a useful way of thinking of such important individuals within our society (see Definition 29).
@@
In the present paper we will focus on heuristic measures, revolving around local centralities due to lower time complexity: while global metrics take into consideration more complete information about the graph, they are often computationally expensive. Furthermore, Hu et al. [30] claimed that global influence can be approximated by local influence from neighbors of order around three to four, which means local measures may be more cost-effective. As mentioned before, one should keep in mind that social networks are constantly changing. In 2004, Liben-Nowell and Kleinberg [37] formalized the link prediction problem for social networks as predicting future edges based on knowledge of the current graph. Hasan and Zaki [28] and Lü and Zhou [44] provided surveys of link prediction results. In our work, we shall focus on local similarity metrics and their quasi-local extensions given by Aziz et al. [2], which are heuristic measures similar to centrality.
@@
In 2019, Ghafouri and Khasteh [23] introduced the idea of using a graph modelling technique to engage in link prediction prior to use of greedy algorithms for influence maximization as a way to account from invisible edges between nodes (i.e. edges not observed when collecting data on the network). Singh and Kailasam [55] utilized a similar approach when predicting sets of influential nodes for the next iteration of the graph through link prediction using the a type of Restricted Boltzmann Machine (RBM). In 2023, Yanchenko et al. [65] discussed the use of deep learning link prediction to understand the outcomes of the future networks, then identify influential nodes through the greedy algorithm and dynamic degree discount. We take a similar approach, but focus on heuristic link prediction measures whereas they consider linear regression, deep learning, matrix factorization, and graph neural networks, and more. We also evaluate many other common centrality metrics and algorithms that, to the best of our knowledge, previous literature has not covered. In an effort to lower time complexity, our graph prediction also makes use of expected value to decrease computation cost of influence spread.
We seeks to fill a very important gap in the literature which would allow the understanding of influencer predictions in evolving social networks; that is, the effects of link prediction on heuristic centrality measures. To the best of our knowledge, link prediction similarity metrics have not yet been used with centrality metrics and $k$ node selection algorithms to approximate the effects of time on a dynamic social network.
@@
We shall seek to quantify influence through simple and complex contagion. As stated by Min & Miguel [46], simple contagion is a model from epidemiology that runs on the assumption that there is a fixed probability of transmission between each pair of vertices. Centola and Macy [12] defined complex contagion as a model for actions or information with some perceived social cost, considering both personal resistance to influence and forces like peer pressure and social affirmation: a person is infected only if sufficiently many of their neighbors are.
@@
In general terms, easily accepted information spreads as simple contagion whereas more controversial or effort-requiring topics may be complex contagion. Examples of some categorizations can be seen in Table 1 below. We shall begin our study in Section 2) by studying link prediction and influencer identification separately on dynamic social networks, and then we shall examine influence through simple and complex contagion models in Section 4.2). By putting the above together we shall present our definition of influencer, detail the specific influence model definitions, and introduce our concrete algorithm, the Social Sphere Model (see Section 3).
@@
Viral Memes Simple Contagion Weng et al. [63]
Nonviral Memes Complex Contagion Weng et al. [63]
Spread of Disease Simple Contagion
Health Behaviors Complex Contagion Campbell and Salath’e [9]
Music Simple Contagion Notarmuzi et al. [49]
TV Shows Simple Contagion Notarmuzi et al. [49]
Political/Societal Controversies Complex Contagion Notarmuzi et al. [49]
Social Movements Complex Contagion Centola [11]
@@
Researchers have noted that the performances of the same set of nodes differ noticeably in the simple and complex contagion (see Figure 1 for an example). For example, Watts and Dodds [61] found that while the former places a great deal of emphasis on influencers, the latter says away from choosing traditional influencers as initial nodes, especially when considering the relative costs (traditionally influential nodes are likely more expensive hires).
@@
In Section 5, and through the use of the Social Sphere Model, we observe that influence generally spreads much faster in the simple contagion model than in the complex. We also notice the following:
  • better performance for some parameters in our modified prediction metrics and algorithms when compared to the originals (see Section 2);
  • high similarity between our model’s predicted influencers and the ‘true’ future influencers in both selection and collective influence, suggesting the effectiveness of our model.
Our paper is organized as follows: in Section 2, we explore previous works around heuristic measures in link prediction and influencer identification. We explain the Social Sphere Model and our mathematical definition for influencers in Section 3. Our methodology is detailed in Section 4, with evaluation methods in Section 4.1, concrete contagion models in Section 4.2, and experimental setup in Section 4.3. We examine our model’s influence in simple and complex contagion through analysis of our algorithm’s results on random graphs in Section 5. Finally, we give a summary and future directions for our research in Section 6.
2 Social Networks
@@
where $w(u,v)$ is the weight of edge $uv$. Such weights may represent a variety of parameters of the social network, and in this paper, we consider weights to be approximations for probabilities of influence transmission, be it information, disease, or more. In particular, we take inspiration from Goyal et al. [24], who took the weights $p(i,j)$ to be the total number of interactions over the total number of trials. In order to study such models, one may assume that the probability of interaction between two nodes during a fixed interval of time is constant, in which case each interval of time can be considered as a Bernoulli trial for temporal datasets.
2.1 Weight functions
@@ 2.2 Link Prediction
  • common neighbors (e.g. [48]),
  • local path (e.g. [70]),
  • Jaccard coefficient (e.g. [37]),
  • resource allocation (e.g. [70]),
  • quasi-local resource allocation (e.g. [2]),
  • RA22-2- 2, and
  • quasi-local RA22-2- 2.
Common Neighbors. Considering collaboration networks, Newman proposed in [48] that the more common neighbors two nodes share, the more likely they are to form a link in the future, leading to the following definition:
@@
Murata and Moriyasu [47] amended common neighbors for weighted networks to be dependent on the sum of the weights between the nodes summed over common neighbors leading to the following weighted definition:
@@
Local Path (Quasi-Local Common Neighbors). Zhou et al. [70] created the local path metric from the observation that common neighbors as a similarity metric often gives the same scores to many nodes, making the actual rankings ambiguous. They extended the computation of score to include paths of three edges, weighting the longer paths by some $\epsilon$:
@@
Here, we take $\epsilon=10^{-3}$. Moreover, Aziz et al. [2] noted that local path is the quasi-local extension of common neighbors. Extending on the previously mentioned work of Murata and Moriyasu [47], Bai et al. [3] developed a weighted index for local path (weights here are adjusted to match the latter):
@@
Jaccard Coefficient. Liben-Nowell and Kleinberg [37] derived the Jaccard coefficient from information retrieval, with the following definition:
@@
Resource Allocation. Zhou et al. [70] came up with the resource allocation metric based on the eponymous process where each individual has resources to be split evenly among its neighbors, defined as follows:
@@
Lü and Zhou [41] made a variant of the above resource allocation for weighted graphs, leading to the following:
@@
Quasi-Local Resource Allocation. Aziz et al. [2] made a quasi-local extension for RA, where one considers paths of three edges in addition to paths of two edges. In other words:
@@
Quasi-Local variant of RA22-2- 2 Inspired by Aziz et al. [2], we also create a quasi-local variant for RA22-2- 2:
@@ 2.3 Node Centrality
Degree Centrality $O(|E|)$[15] $O(|E|)$[15] local [59]
Coreness $O(|E|)$[4] $O(|E|)$ iterative [59]
H-index $O(|V|\langle k\rangle)$ [40] $O(|V|\langle k\rangle)$ local [59]
LocalRank $O(|V|\langle k\rangle^{2})$ [14] N/A local [59]
Clustering Coefficient (and ClusterRank) $O(|V|\max_{v\in V}k_{v}^{2})$ [51] $O(|V|\max_{v\in V}k_{v}^{2})$ [59]
Closeness $O(|V||E|)$ [60] $O(|V||E|+|V|^{2}\log|V|)$ [20] path-based [42], global [59]
Betweenness $O(|V||E|)$ [7] $O(|V||E|+|V|^{2}\log|V|)$ [7] path-based [42], global [59]
Eigenvector $O(|V|^{2}\cdot\text{iteration \#})$ [53] $O(|V|^{2}\cdot\text{iteration \#})$ [53] global, iterative [59]
PageRank $O(|E|\cdot\text{iteration \#})$ [58] $O(|E|\cdot\text{iteration \#})$ [58] iterative [59]
LeaderRank $O(|E|\cdot\text{iteration \#})$ $O(|E|\cdot\text{iteration \#})$ iterative [59]
Balanced Index $O(\langle k\rangle V+V\log V)$ [31] $O(\langle k\rangle V+V\log V)$ local
@@
Vital node identification is a crucial part of social network analysis, applicable to many field ranging from marketing to epidemiology. Centrality methods, some of the most common of these processes, offer a window into understanding influence dynamics within social networks. Mathematically, centrality methods are based on functions that assign a real-number score to each $v\in V$, essentially ranking the nodes in terms of influence [42].
@@
  • degree;
  • coreness, introduced by Seidman [52];
  • H-index, introduced by Hirsch [29];
  • LocalRank, introduced by Chen et al. [14];
  • ClusterRank, introduced by Chen et al. [13];
  • closeness, introduced by Bavelas [5];
  • betweenness, introduced by Freeman [21];
  • eigenvector, introduced by Landau [35];
  • PageRank, introduced by Brin and Page [8];
  • LeaderRank, introduced by Lü et al. [43];
  • balanced index, introduced by Karampourniotis et al. [31], and
  • complex path, introduced by Guilbeault and Centola [25].
@@
Coreness. In 1983, Seidman [52] defined the $k$-core as a measure of the network cohesion of a node, where a subgraph is a $k$-core if it has minimum degree at least $k$. Batagelj and Zaveršnik [4] define a $k$-core decomposition algorithm in Algorithm 1.
Algorithm 1 $k$-shell decomposition [4]
@@
Later on, Kitsak et al. [33] took the core numbers for each vertex as its corenness value. Eidsaa and Almaas [18] created a weighted version of $k$-core, where node strength is used rather than degree in the decomposition and when a node $u$ is selected (has the minimum strength in the pruned graph), it is in the $\lceil s_{u}\rceil$-core. A similar definition can be taken using $\lfloor s_{u}\rfloor$, and the algorithm is modified accordingly in Algorithm 2 below.
@@
H-index. Along the same lines of study, but within the academic world, Hirsch [29] proposed the H-index to measure the impact of a scientist’s research outputs through considering citations per paper. Lü et al. [45] then used it in the context of centrality metrics, introducing the following definition:
@@
From the above perspective, the classical H-index can be seen as the $1$-order H-index. Lü et al. [45] proved that the H-index forms a middle ground between a progression from degree centrality to coreness, where the 00-order index is degree and the $\infty$-order index converges to the $k$-core score. Finally, it is interesting to note that Zhao et al. [68] defined the w-lobby index, a metric for weighted networks similar to the H-index, as follows:
@@
LocalRank. Aiming to find a more effective local method due to the high time complexities of global and path-based methods, Chen et al. [14] introduced the following extension of degree centrality to more than just first-order neighborhoods:
@@
ClusterRank. Watts and Strogatz [62] defined the clustering coefficient around the percent of existing links between a node’s neighbors over the total potential amount to quantify how clustered the network is:
@@
Working with the intuition that nodes in more clustered parts of the network may be more redundant and thus have less influence, Chen et al. [13] then used the clustering coefficients of nodes to define the following centrality metric:
@@
Closeness. Bavelas introduced in 1948 [5] the idea of closeness centrality as proportional to the reciprocal of the sum of the shortest path lengths from each pair of vertices $u\neq v$. Rephrasing, Freeman [22] defines the concept of closeness centrality as a measure of the relative proximity of a node $p_{k}$ to the rest of the graph $G$, defined as
@@
In a survey of influencer identification metrics, Lü et al. [42] note that disconnected networks then pose a problem as distances go to infinity, making the nodes indistinguishable by closeness centrality (as they are all then 00). They then give the following modification:
@@
Betweenness. Betweenness centrality considers how critical a node $i$ is to the spread of information, e.g. what percentage of shortest paths going from one node to another must pass through $i$. First formally defined it in 1977 by Freeman [21], it can be understood along with an efficient algorithm for computing betweenness centrality, through the work of Brandes [7]:
@@
Eigenvector. In [35] Landau gave a clarified overview of a previous paper of his detailing the use of eigenvectors to rank chess players, which has been credited with being the first use of eigenvector centrality. Later on, Lü et al. [42] defined the eigenvector centrality scores $EC$ of $G$ as the vector $\vec{x}$ with components $x_{i}=EC_{i}$ such that $\vec{x}:=\frac{1}{\lambda}A\vec{x}$ for the largest eigenvector $\lambda$ of $A$. In other words,
@@
PageRank. PageRank is a search algorithm presented by Brin and Page [8] in a paper where they introduced a Google prototype, modeled on the behavior of a bored, link-clicking user. They defined the algorithm as follows:
@@
Xing and Ghorbani [64] modified the PageRank algorithm for weighted graphs:
@@
LeaderRank. Inspired by PageRank, Lü et al. [43] created the LeaderRank algorithm. Their intention was to include a “ground node" connected to every node in the network, called node $n+1$, and then used the following recursive relationship to solve for their influences:
@@
Karampourniotis et al. [31] introduced the Balanced Index metric for complex contagion by taking a weighted combination the metrics of resistance, degree, and neighbors’ degrees. In the context of threshold models, the resistance $r_{v}$ of a node $v$ is defined as a measure of the current threshold of each node, where uninfected nodes $v$ initially begin with $r_{v}=\theta_{v}\cdot s(v)$ and experience decreasing resistance as more of its neighbors become infected. At resistance 00 the node also becomes infected. The neighbors’ degree comes from Karampourniotis et al.’s [31] indirect drop of resistance metric, which for node $v$ takes the sum over the nodes that $v$ could infect in a single turn (i.e. the nodes $j$ with remaining resistance less than $\frac{w(i,j)}{s(i)}$). The balanced index metric is thus as follows:
@@
Complex Path Centrality. Seeking to extend the path length measures traditionally used in simple contagion to the wide bridges (reinforcing ties) used in complex contagion paths, Guilbeault and Centola [25] created complex path centrality, based on the following: